# Locally Repairable Convertible Codes: Improved Lower Bound and General Construction

### v1

@@ Abstract
In this paper, we consider the convertible code with locally repairable property. We present an improved lower bound on access cost associated with $(r,\delta)$. Then, we provide a general construction of convertible codes with optimal access cost which shows that those codes can be with super-linear length or maximum repairable property. Additionally, employing the known locally repairable codes with super-linear length or maximum repairable property, we provide explicit constructions of convertible codes with super-linear length or maximum repairable property.
Index Terms:
Convertible codes, locally repairable codes, access cost, super-linear length, maximally recoverable codes.
@@ I Introduction
With the rapid expansion of data in large-scale cloud storage and distributed file systems such as Windows Azure Storage and Google File System (GoogleFS), disk failures have become the norm rather than an exception. To against such failures, the simplest approach is to replicate data packets across multiple disks. However, this method incurs significant storage overhead. As an alternative, coding techniques have been developed to enhance storage efficiency.
Considering various aspects of distributed storage system performance, researchers have introduced a range of coding techniques. For instance, to maximize failure tolerance while minimizing redundancy, they have explored maximum distance separable (MDS) codes. To reduce the data transmitted during the repair process, regenerating codes with optimal repair bandwidth have been proposed [10]. To limit the number of nodes involved in data recovery, locally repairable codes have been designed [12]. Over the past decade, significant progress has been made in coding techniques for distributed storage systems based on these performance criteria. For further details, see: Codes with optimal repair bandwidth [10, 14, 45, 49, 44, 27, 25, 43, 50, 29, 17, 16, 8, 46]; Optimal locally repairable codes [18, 12, 37, 41, 1, 22, 28, 48, 5, 6, 7, 15, 39]; Codes with good availability [38, 42, 2, 4, 20]; Codes with optimal update and access properties [43, 35, 26, 9] and references therein.
All these coding schemes are designed under a fixed code rate, which corresponding to a preset failure rate for the storage system. However, device failure rates are inherently dynamic. To solve this problem, we usually preset a low code rate corresponding a worst device failure rate, which may cause waste during the device stateable period. To overcome this drawback, Kadekodi et al. [21] introduced a cost-effective framework that dynamically adapts code rates and redundancy levels to fluctuating device failure rates, thereby optimizing storage efficiency. This process, termed code conversion [32], requires transitioning encoded data from an initial code $[n_{I},k_{I}]$ to a final code $[n_{F},k_{F}]$ with distinct parameters. Code conversion operates in two primary modes:
@@
The traditional code conversion method involves re-encoding the original data from an initial code $\mathcal{C}_{I}$ to the final code $\mathcal{C}_{F}$. However, this approach requires retrieving a vast number of symbols, decoding the entire dataset, transmitting the data across the network, and subsequently re-encoding it. This results in considerable overhead in terms of network bandwidth, I/O load, and CPU resource utilization within the cluster, thereby impeding the system’s efficiency. To address these limitations, Maturana and Rashmi [32] pioneered the theory of convertible codes enabling parameter transitions without full data reconstruction. Building upon their framework, subsequent studies [33, 34, 31] established fundamental theorems for maximum distance separable (MDS) code conversions. Subsequently, [23] and [11] provided a class of Reed-Solomon (RS) codes with optimal read and write access costs based on a construction of the locally repairable codes (LRCs) [41] and the extended generalized RS codes [19], respectively,
@@
Analogous to MDS codes, storage systems employing LRCs confront the inefficiency of employing code with a fixed rate, but the failure rate of storage devices will change over time. Consequently, the design of LRCs that facilitate efficient code conversions presents both practical and theoretical significance. Maturana and Rashmi [34] first proposed a new construction technique for designing LRCs that can perform code conversion at a lower cost than the default approach by direct computation, projection, piggybacks, projected piggybacks, linear combination, and instance reassignment. Subsequently, Kong [23] established a lower bound on the access cost of code conversion for general LRCs and proposed constructions of LRCs that can perform code conversions with access cost achieving this bound with equality based on LRCs first introduced in [41].
For locally repairable codes, existing results suggest that the code length can grow super-linearly with respect to the alphabet size for some parameter regime, or they allow for recoverable erasure patterns beyond the minimum Hamming distance. In contrast, for convertible codes, known results impose a linear relationship between the alphabet size and the code length while restricting recoverable erasure patterns to sizes smaller than the minimum Hamming distance. This naturally raises the question: Can locally repairable convertible codes achieve super-linear length or support recoverable erasure patterns beyond the minimum Hamming distance? To address this, we first establish an improved lower bound on the access cost associated with $(r,\delta)$ locality. We then present a general construction of convertible codes with optimal access cost, demonstrating that such codes can indeed achieve super-linear length or support recoverable erasure patterns beyond the minimum Hamming distance. Finally, as concrete examples, we construct explicit convertible codes with super-linear length by leveraging known locally repairable codes with the same property. In Table I and Table II, we summarize and compare known related results with our results.
@@
The subsequent sections of the paper are organized as follows: Section II reviews notation and preliminaries. Section III presents improved lower bounds on access cost for locally repairable convertible codes in the merge regime. Section IV proposes a general construction of access-optimal locally repairable convertible codes. Section V proposes explicit access-optimal locally repairable convertible codes with super-linear length or maximally repairable property. Finally, Section VI concludes this paper.
@@
$\mathcal{C}_{I}$: $[n_{I},k_{I}]_{q}$ code $\mathcal{C}_{F}$: $[n_{F},k_{F},d]_{q}$-LRC with $(r,\delta)_{a}$-locality $\rho_{r}\geqslant\begin{cases}k_{F},\mbox{if }\Delta\leqslant 0\text{ or }d>n_{I}-k_{I}+1,\\ t\left(k_{I}-\lceil\frac{r\Delta}{r+1}\rceil\right),\mbox{otherwise},\end{cases}$ where $\Delta=n_{F}-2d-(t-1)k_{I}+3-(\lceil\frac{(t-1)k_{I}}{r}\rceil)$ $\rho_{w}\geqslant n_{F}-t(\Delta+d-1)$ [23, Theorem IV.4]
$\mathcal{C}_{I}$: $[n_{I},k_{I}]_{q}$ code $\mathcal{C}_{F}$: $[n_{F},k_{F},d]_{q}$-LRC with $(r,\delta)_{a}$-locality $\rho_{r}\geqslant\begin{cases}k_{F},\mbox{if }\Delta\leqslant 0\text{ or }d>n_{I}-k_{I}+1,\\ t\left(k_{I}-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor\right),\mbox{otherwise},\end{cases}$ where $\Delta=n_{F}-2d-(t-1)k_{I}+2-(\lceil\frac{(t-1)k_{I}}{r}\rceil-1)(\delta-1)$ $\rho_{w}\geqslant n_{F}-t(\Delta+d-1)$ Theorem 2
TABLE II: Known Constructions of $(t,1)_{q}$ LRCCs within Merge Regime with Parameters $(n_{I},k_{I},r_{I},\delta_{I};n_{F},k_{F}=tk_{I},r_{F},\delta_{F})_{a}$
@@
Access optimal LRCCs [23, Construction III,IV] $r_{I}=r_{F}=r$ $\delta_{I}=\delta_{F}=2$ $r|k_{I},r|k_{F}$ $(r+1)|n_{I},(r+1)|n_{F}$ $\frac{k_{I}(r+1)}{r}~{}|~{}(q-1)$ and $q\geqslant\frac{k_{I}(r+1)}{r}\max\{t+1,\lceil\frac{rn_{I}}{(r+1)k_{I}}\rceil+1\}+1$
General access optimal LRCCs (Theorem 3) $t|\ell,\ell/t>g-\ell\geqslant h\geqslant 0$ $r_{I}=r_{F}=r$ $\delta_{I}=\delta_{F}=\delta\geqslant 2$ $n_{I}=(\ell/t+g-\ell)(r+\delta-1)$ $n_{F}=(\ell+h)(r+\delta-1)$ $k_{I}=\ell r/t,k_{F}=\ell r$ Determined by the field size requirement of $\mathcal{C}_{F}$ Examples: super-linear length (Corollary 2) maximum recoverable (Corollary 4)
LRCs with information locality [47] $\mathcal{C}_{I}$: $[20,12]$-LRC with $(2,2)$-locality $\mathcal{C}_{F}$: $[16,12]$-LRC with $(6,2)$-locality, or vice versa (In this box, we show the access cost.) $6$(read) $+2$(write), $\mathcal{C}_{I}\to\mathcal{C}_{F}$, $10$(read) $+6$(write), $\mathcal{C}_{F}\to\mathcal{C}_{I}$. changed locality
@@ II Preliminary
  • An $[n,k,d]_{q}$ linear code $\mathcal{C}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with the minimal Hamming distance $d$;
  • For a set $S\subseteq[n]$, let $\mathcal{C}|_{S}$ denote the punctured code formed by all the codewords of $\mathcal{C}$ that delete symbols $c_{j}$ for all $j\in[n]\setminus S$;
  • Let $O_{m\times n}$ and $I_{m\times n}$ denote the $m\times n$ zero matrix and identity matrix, respectively. If $m,n$ are implicit from context, we simply write $O$ and $I$ for short;
@@
    $H=\begin{pmatrix}A_{1}&O&\cdots&O\\ O&A_{2}&\cdots&O\\ \vdots&\vdots&\ddots&\vdots\\ O&O&\cdots&A_{g}\\ B_{1}&B_{2}&\cdots&B_{g}\end{pmatrix},$ (1)
@@
    $H||_{\{i_{1},i_{2},\dots,i_{t}\}}\triangleq\begin{pmatrix}A_{i_{1}}&O&\cdots&O\\ O&A_{i_{2}}&\cdots&O\\ \vdots&\vdots&\ddots&\vdots\\ O&O&\cdots&A_{i_{t}}\\ B_{i_{1}}&B_{i_{2}}&\cdots&B_{i_{t}}\end{pmatrix},$ (2)
    where $i_{1}<i_{2}<\dots<i_{t}$ and $i_{1},i_{2},\dots,i_{t}\in[g]$;
@@
It is well known that the following lemma characterizes the distance property.
Lemma 1 ([19, Theorem 1.5.7]):
Let $\mathcal{C}$ be an $[n,k,d]_{q}$ code. For any set $S\subseteq[n]$, if $n-|S|<d$, then the punctured code $\mathcal{C}|_{S}$ has dimension $k$.
@@ II-A Locally Repairable Codes
Let $\mathcal{C}$ be an $[n,k,d]_{q}$ code. The $i$th code symbol $c_{i}$ of $\mathcal{C}$ is called to have $(r,\delta)$-locality if there exists a subset $S_{i}\subset[n]$ (an $(r,\delta)$-repair set) satisfying
@@
For an $[n,k,d]_{q}$ linear code $\mathcal{C}$, if all the $n$ code symbols have $(r,\delta)$-locality, then it is said to have all symbol $(r,\delta)$-locality, denoted as locally repair codes (LRCs) with $(r,\delta)_{a}$-locality. For a LRC with $(r,\delta)_{a}$-locality, Lemma 1 and Definition 1 imply that there exists a subset $\mathrm{Loc}(i)\subseteq S_{i}$ with $|\mathrm{Loc}(i)|\leqslant r$ and $i\not\in\mathrm{Loc}(i)$ such that
@@
where $\lambda_{j}\in\mathbb{F}_{q}^{*}$ for all $j\in\mathrm{Loc}(i)$. Note that, for given $i\in[n]$, $\mathrm{Loc}(i)$ may be not unique. In the following, we use the notation $\mathrm{Loc}(i)$ to denote one of the sets with the above property for the $i$th code symbol $c_{i}$.
An upper bound on the minimum Hamming distance of linear codes with $(r,\delta)_{a}$-locality was derived as follows.
@@
The minimum distance of an $[n,k,d]_{q}$ code $\mathcal{C}$ with $(r,\delta)_{a}$-locality is upper bounded by
@@
An LRC code is called optimal if the upper bound (3) is achieved. The repair sets of linear code $\mathcal{C}$ with $(r,\delta)_{a}$-locality has the following important property.
Lemma 3 ([40, Theorem 9]):
Let $\mathcal{C}$ be an optimal linear LRC with $(r,\delta)_{a}$-locality. If $r\mid k$ and $r<k$, then there are $t$ repair sets $S_{i_{1}},\dots,S_{i_{t}}$ are mutually disjoint where $S_{i_{j}}\subseteq[n]$ for $j\in[t]$, and $|S_{i_{j}}|=r+\delta-1$ for all $j\in[t]$, and the punctured code $\mathcal{C}|_{S_{i_{j}}}$ is an $[r+\delta-1,r,\delta]$ MDS code. In particular, we have $(r+\delta-1)\mid n$, and $t=\frac{n}{r+\delta-1}$.

### v2

@@ Abstract
In this paper, we consider convertible codes with the locally repairable property. We present an improved lower bound on access cost associated with $(r,\delta\geqslant 2)$-locality, which extends the known lower bound only related to $r$. We then provide a general construction of convertible codes with optimal access cost which shows that these codes can have super-linear length or maximum repairable property. Furthermore, we propose explicit constructions of convertible codes with the final code achieving super-linear length or maximum repairable property when the initial codes have super-linear length or maximum repairable property respectively. Specially, compared with known constructions, our construction is applicable to the case $\delta>2$ for the first time.
@@ I Introduction
The proliferation of large-scale cloud storage systems and distributed file architectures (e.g., Windows Azure Storage, Google File System) has transformed disk failures from exceptional events into commonplace occurrences. To combat these failures, conventional replication strategies create multiple copies of data blocks across storage nodes, albeit at substantial storage costs. This limitation has driven the development of erasure coding techniques that achieve comparable reliability with significantly improved storage efficiency.
Considering various aspects of distributed storage system performance, researchers have introduced a range of coding techniques. For instance, to maximize failure tolerance while minimizing redundancy, they have explored maximum distance separable (MDS) codes. On the one hand, to reduce the data transmitted during the repair process, regenerating codes with optimal repair bandwidth have been proposed [12]. On the other hand, to limit the number of nodes involved in data recovery, locally repairable codes (LRCs) have been designed [15]. Over the past decade, significant progress has been made in coding techniques for distributed storage systems based on these performance criteria. For further details, refer to the following: Codes with optimal repair bandwidth [12, 17, 51, 55, 50, 32, 30, 49, 56, 34, 21, 20, 10, 52, 29]; Optimal locally repairable codes [22, 15, 42, 47, 2, 26, 33, 54, 6, 7, 9, 19, 44]; Codes with good availability [43, 48, 3, 5, 24]; Codes with optimal update and access properties [49, 40, 31, 11] and references therein.
All these coding schemes are designed under a fixed code rate, which corresponds to a preset failure rate for the storage system. However, device failure rates are inherently dynamic. Thus, we usually preset a low code rate corresponding to the worst device failure rate, which may cause inefficiency during the device’s stable period. To overcome this drawback, Kadekodi et al. [25] introduced a cost-effective framework that can dynamically adapt code rates and redundancy levels to the fluctuations in the device failure rate, thereby optimizing storage efficiency. This process, termed code conversion [37], requires transitioning encoded data from an initial code $n_{I},k_{I}]$ to a final code $n_{F},k_{F}]$ with distinct parameters. Code conversion operates in two primary modes:
@@
The traditional code conversion method involves re-encoding the original data from an initial code $\mathcal{C}_{I}$ to the final code $\mathcal{C}_{F}$. Nevertheless, this approach requires retrieving a vast number of symbols, decoding the entire dataset, transmitting the data across the network, and subsequently re-encoding it. This results in considerable overhead in terms of network bandwidth, I/O load, and CPU resource utilization within the cluster, hence impeding the system’s efficiency. To address these limitations, Maturana and Rashmi [37] pioneered the theory of convertible codes enabling parameter transitions without full data reconstruction. Building upon their framework, the studies [38, 39, 36] established fundamental theories for MDS code conversions. Subsequently, [27] and [13] provided a class of convertible codes with optimal read and write access costs based on RS codes and the extended generalized RS codes [23], respectively.
@@
Analogous to MDS codes, storage systems employing LRCs face the inefficiency of using a code with a fixed rate, but the failure rate of storage devices changes over time. Consequently, the design of LRCs that facilitate efficient code conversions is of both practical and theoretical significance. Xia et al. [53] first proposed a conversion method from the initial (2, 2)-locality to the final (6, 2)-locality while minimizing the number of node read. Subsequent constructions mainly considered the same locality between the initial and final codes. In [39], Maturana and Rashmi proposed a new construction technique for designing locally repairable convertible codes (LRCCs) that can perform code conversion at a lower cost than the default approach by direct computation, projection, piggybacks, projected piggybacks, linear combination, and instance reassignment. Recently, Kong [27] established a lower bound on the access cost of LRCCs and proposed constructions of LRCCs that can perform code conversions with access cost achieving this bound with equality based on LRCs introduced in [47].
The long-standing MDS conjecture asserts that no non-trivial MDS codes (with distance $d>2$) exist beyond length $q+1$ over an alphabet of size $q$, with precisely two exceptional cases: when $q$ is even with $k=3$, or k=q1k=q-1, the maximal allowable length extends to $q+2$. This conjecture was resolved for prime fields by Ball [1]. For MDS convertible codes, Ge et al., [13] presented constructions with field size $q=n_{F}-1$, where $n_{F}$ is the code length of the final code. The situation differs fundamentally for optimal LRCs. The studies [18, 6, 28] proposed explicit constructions of optimal LRCs with super-linear length, i.e., $n=O(q)$. This naturally raises the first open question [27]:
Question 1: Can locally repairable convertible codes satisfy that the final code preserves super-linear length when the initial codes have super-linear length?
While the benefits of locality are well-established, this feature entails a critical trade-off between the code rates and minimum Hamming distances. The maximally recoverable (MR) codes [15] can correct erasures beyond their guaranteed minimum distance. As a specialized subclass of LRCs, MR codes can correct all erasure patterns that are information-theoretically correctible given the prescribed dependency relations between data symbols and parity symbols. This naturally raises the second question:
Question 2: Can locally repairable convertible codes satisfy that the final code preserves maximally recoverable property when the initial codes inherently exhibit maximally recoverable property?
To address these two questions, we first establish an improved lower bound on the access cost for convertible codes with $(r,\delta)$-locality. A comprehensive comparison with existing bounds is presented in Table I. We then propose the existence conditions and a general construction for access-optimal locally repairable convertible codes with a special class of parameters. Finally, as concrete examples, we construct explicit convertible codes with super-linear length (i.e., $n_{F}=\Omega(q^{\delta})$ and $n_{I}=\Omega_{t}(n_{F}/t)$, where $n_{F}$ and $n_{I}$ are the code lengths of the final code and initial codes respectively, and $t$ is the number of initial merged codewords), or maximally recoverable property (i.e., both the final and initial codes are MR codes), based on the general construction. To the best of our knowledge, these are the first explicit constructions of LRCCs with super-linear length or maximally recoverable property. In Table II, we summarize and compare known constructions with ours which are the first constructions with $\delta>2$.
@@
The remainder sections of the paper are organized as follows: Section II reviews notation and preliminaries. Section III presents improved lower bounds on access cost for locally repairable convertible codes in the merge regime. Section IV proposes a general construction of access-optimal locally repairable convertible codes. Section V proposes explicit access-optimal locally repairable convertible codes with super-linear length or maximally repairable property. Finally, Section VI concludes this paper.
@@
$\mathcal{C}_{I}$: $n_{I},k_{I}]_{q}$ code
$\mathcal{C}_{F}$: $n_{F},k_{F},d]_{q}-$LRC
with $(r,\delta)_{a}$-locality
$\rho_{r}\geqslant\begin{cases}k_{F},\mbox{if }\overline{\Delta}\leqslant 0\text{ or }d>n_{I}-k_{I}+1,\\ t\left(k_{I}-\lceil\frac{r\overline{\Delta}}{r+1}\rceil\right),\mbox{otherwise},\end{cases}$
where $\overline{\Delta}=n_{F}-2d-(t-1)k_{I}+3-\lceil\frac{(t-1)k_{I}}{r}\rceil$
$\rho_{w}\geqslant n_{F}-t(\overline{\Delta}+d-1)$ [27, Theorem IV.4]
$\mathcal{C}_{I}$: $n_{I},k_{I}]_{q}$ code
$\mathcal{C}_{F}$: $n_{F},k_{F},d]_{q}-$LRC
with $(r,\delta)_{a}$-locality
$\rho_{r}\geqslant\begin{cases}k_{F},\mbox{if }\Delta\leqslant 0\text{ or }d>n_{I}-k_{I}+1,\\ t\left(k_{I}-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right),\mbox{otherwise},\end{cases}$
where $\Delta=n_{F}-2d-(t-1)k_{I}+2-(\lceil\frac{(t-1)k_{I}}{r}\rceil-1)(\delta-1)$
$\rho_{w}\geqslant n_{F}-t(\Delta+d-1)$ Corollary 1
TABLE II: Known Constructions of $(t,1)_{q}$ Locally Repairable Convertible Codes within Merge Regime with Parameters $(n_{I},k_{I},r_{I},\delta_{I};n_{F},k_{F}=tk_{I},r_{F},\delta_{F})_{a}$
@@
Access optimal LRCCs
[27, Construction III,IV]
$r_{I}=r_{F}=r$
$\delta_{I}=\delta_{F}=2$
$r|k_{I},r|k_{F}$
$(r+1)|n_{I},(r+1)|n_{F}$
$\frac{k_{I}(r+1)}{r}~|~(q-1)$ and
$q\geqslant\frac{k_{I}(r+1)}{r}\max\{t+1,\lceil\frac{rn_{I}}{(r+1)k_{I}}\rceil+1\}+1$
General access optimal LRCCs
(Construction A)
$t|\ell,\ell/t>g-\ell\geqslant h\geqslant 0$
$r_{I}=r_{F}=r$
$\delta_{I}=\delta_{F}=\delta\geqslant 2$
$n_{I}=(\ell/t+g-\ell)(r+\delta-1)$
$n_{F}=(\ell+h)(r+\delta-1)$
$k_{I}=\ell r/t,k_{F}=\ell r$
Determined by the field size
requirement of $\mathcal{C}_{I}$ and $\mathcal{C}_{F}$
super-linear length
(Theorem 4)
or maximally recoverable
(Theorem 5)
LRCCs with information locality
[53]
$t=1$,
$r_{I}=2,r_{F}=6$
$\delta_{I}=\delta_{F}=2$
$n_{I}=20,n_{F}=16$
$k_{I}=k_{F}=12$
$q=2^{4}$ changed locality
@@ II Preliminary
  • An $n,k,d]_{q}$ linear code $\mathcal{C}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with the minimal Hamming distance $d$;
  • For a set $S\subseteq[n]$, let $\mathcal{C}|_{S}$ denote the punctured code formed by deleting symbols $c_{j}$ for all $j\in[n]\setminus S$ from all codewords of $\mathcal{C}$;
  • Let $\bm{0}_{m\times n}$ and $I_{m\times n}$ denote the $m\times n$ zero matrix and identity matrix, respectively. If $m,n$ are clear from context, we simply write $\bm{0}$ and $I$ for short;
@@
    $H=\begin{pmatrix}A_{1}&\bm{0}&\cdots&\bm{0}\\ \bm{0}&A_{2}&\cdots&\bm{0}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{0}&\bm{0}&\cdots&A_{g}\\ B_{1}&B_{2}&\cdots&B_{g}\end{pmatrix}$, (1)
@@
    $H||_{\{i_{1},i_{2},\dots,i_{t}\}}\triangleq\begin{pmatrix}A_{i_{1}}&\bm{0}&\cdots&\bm{0}\\ \bm{0}&A_{i_{2}}&\cdots&\bm{0}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{0}&\bm{0}&\cdots&A_{i_{t}}\\ B_{i_{1}}&B_{i_{2}}&\cdots&B_{i_{t}}\end{pmatrix}$, (2)
    where $1\leqslant i_{1}<i_{2}<\dots<i_{t}\leqslant g$;
@@
It is well known that the following lemma characterizes the distance property of a punctured code.
Lemma 1 ([23, Theorem 1.5.7]):
Let $\mathcal{C}$ be an $n,k,d]_{q}$ code. For any set $S\subseteq[n]$, if n|S|<dn-|S|<d, then the punctured code $\mathcal{C}|_{S}$ has dimension $k$.
@@ II-A Locally Repairable Codes
Let $\mathcal{C}$ be an $n,k,d]_{q}$ code. The $i$th code symbol $c_{i}$ of $\mathcal{C}$ is said to have $(r,\delta)$-locality if there exists a subset $S_{i}\subset[n]$ (an $(r,\delta)$-repair set) satisfying
@@
For an $n,k,d]_{q}$ linear code $\mathcal{C}$, if all the $n$ code symbols have $(r,\delta)$-locality, then it is said to have all symbol $(r,\delta)$-locality, denoted as locally repairable codes (LRCs) with $(r,\delta)_{a}$-locality. For an LRC with $(r,\delta)_{a}$-locality, Lemma 1 and Definition 1 imply that there exists a subset $\mathrm{Loc}(i)\subseteq S_{i}$ with $|\mathrm{Loc}(i)|\leqslant r$ and $i\not\in\mathrm{Loc}(i)$, such that
@@
where $\lambda_{j}\in\mathbb{F}_{q}^{*}$ for all $j\in\mathrm{Loc}(i)$. Note that, for given $i\in[n]$, $\mathrm{Loc}(i)$ may be not unique. In the following, we use the notation $\mathrm{Loc}(i)$ to denote one of such sets for the $i$th code symbol $c_{i}$.
Some upper bounds on the minimum Hamming distance of linear codes with $(r,\delta)_{a}$-locality were derived as follows.
@@
The minimum distance of an $n,k,d]_{q}$ code $\mathcal{C}$ with $(r,\delta)_{a}$-locality is upper bounded by
@@
Lemma 3 ([28, Remark II.4]):
$d\leqslant n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1).$ (4)
Given $n,k,r,\delta,q$, let $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)$ is the largest possible minimum Hamming distance of an $n$-length LRCs with $(r,\delta)_{a}$-locality for given finite field $\mathbb{F}_{q}$. When the bound in Lemma 2 or Lemma 3 is tight, then $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left(\left\lceil\frac{k}{r}\right\rceil-1\right)(\delta-1)$ or $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1)$. An $n,k,d]_{q}$ LRC with $(r,\delta)_{a}$-locality is called optimal if $d=d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)$. To the convenience of analysis, we assume that
$d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\phi(k,r,\delta),$ (5)
where $\phi(k,r,\delta)$ is determined by $k,r,\delta$.
The repair sets of linear code $\mathcal{C}$ with $(r,\delta)_{a}$-locality have the following important property.
Lemma 4 ([6, Lemma 2]):
Let $\mathcal{C}$ be an optimal linear LRC with $(r,\delta)_{a}$-locality and achieving the minimum Hamming distance bound in (3). If $r\mid k$ and $r<k$, then there are $t$ repair sets $S_{i_{1}},\dots,S_{i_{t}}$ are mutually disjoint where $S_{i_{j}}\subseteq[n]$ for $j\in[t]$, and $|S_{i_{j}}|=r+\delta-1$ for all $j\in[t]$, and the punctured code $\mathcal{C}|_{S_{i_{j}}}$ is an $r+\delta-1,r,\delta]$ MDS code. In particular, we have $(r+\delta-1)\mid n$, and $t=\frac{n}{r+\delta-1}$. Furthermore, the code $\mathcal{C}$ has a parity check matrix $H$ of the form (1) where $g=t$ and $A_{j}$ is the parity check matrix of $\mathcal{C}|_{S_{i_{j}}}$ for $j\in[t]$ up to a rearrangement of the code coordinates.